OpenStudy (anonymous):

Rewrite the equation in Ax+By=C This is what I have so far: y-3=2(x-5) y-2=2x-10

jimthompson5910 (jim_thompson5910):

since we want it in the form Ax+By=C, you basically have to get the x & y terms on the left side get everything else to the right side

OpenStudy (anonymous):

So subtract 2x. 2x-y-3=-10

OpenStudy (anonymous):

I mean -2x+y-3=-10

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

Then -2x+y=-7

jimthompson5910 (jim_thompson5910):

good, y-3=2(x-5) y-3=2x-10 -2x+y = -10+3 -2x+y = -7

jimthompson5910 (jim_thompson5910):

now some books require A to be positive to do that, we multiply both sides by -1 -2x+y = -7 -1*(-2x+y) = -1*(-7) 2x-y = 7

OpenStudy (anonymous):

Then its 2x+y=7?

jimthompson5910 (jim_thompson5910):

this step is optional

OpenStudy (anonymous):

ohh

jimthompson5910 (jim_thompson5910):

you forgot to multiply -1 by the +y term to get -y

OpenStudy (anonymous):

Thanks for the clarification. 2x-y=7??

jimthompson5910 (jim_thompson5910):

yep, assuming the requirement your book/teacher needs A > 0

OpenStudy (anonymous):

Yes! Alright thanks for that!

jimthompson5910 (jim_thompson5910):

no problem